# A modern perspective on rational homotopy theory

### v1

@@ 5.2. Function complexes, tensors, and cotensors of $1$-reduced simplicial sets
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
5.3. An exponential law
The simplicial model structure on ${\mathcal{SS}\mathrm{ets}}$ is underpinned by the exponential law for simplicial sets [GJ99, Chapter I, Proposition 5.1]. We formulate an exponential law for pointed simplicial sets for our simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$. For two pointed simplicial sets $X$ and $Y$, the evaluation
$\operatorname{ev}\colon X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\rightarrow Y$
is defined at the level of $n$-simplices by
$\operatorname{ev}_{n}([x,f])=f_{n}([x,\sigma_{n}]),$
where $\sigma_{n}$ is the unique non-degenerate $n$-simplex of $\Delta[n]$. We establish the following exponential law.
Proposition 5.4.
For pointed simplicial sets $X$, $Y$, and $Z$, the map
$\operatorname{ev}_{\ast}\colon\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\rightarrow\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z,Y\right)$
sending $f\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$ to the composite
$X\wedge Z\xrightarrow{\operatorname{id}_{X}\wedge f}X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\xrightarrow{\operatorname{ev}}Y$
is a bijection that is natural in $X$, $Y$, and $Z$.
Remark 5.5.
This natural bijection amounts to the adjunction $X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-)$.
Proof.
The inverse of $\operatorname{ev}_{\ast}$ sends a map $g\colon X\wedge Z\rightarrow Y$ to the map
$\hat{g}\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$
defined by sending an $n$-simplex $z$ of $Z$, which we view as a map $z\colon\Delta[n]\rightarrow Z$, to the composite
[…]
where, since $Z$ is pointed, we use Lemma 2.8 to induce the pointed map $\widetilde{z}\colon\Delta[n]_{+}\rightarrow Z$. ∎
In particular, we use the following consequence of our exponential law.
Corollary 5.6.
For a $1$-reduced simplicial set $X$, there is an adjunction
$X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-).$
Proof.
By Lemma 2.15, the functor $X\wedge-$ maps into ${\mathcal{SS}\mathrm{ets}}_{1}$. The adjunction now follows from Remark 5.5. ∎
5.4. Verification of the adjointness axiom
We use our exponential law to verify that the structure defined in Section 5.2 satisfies the adjointness axiom. We need an auxiliary lemma that is straightforward to check.
Lemma 5.7.
Given two simplicial sets $Z$ and $W$, there is an isomorphism of pointed simplicial sets
$\left(Z_{+}\right)\wedge\left(W_{+}\right)\cong\left(Z\times W\right)_{+}$
that is natural in $Z$ and $W$.
We are ready to verify the adjointness axiom for ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@ 5.4. Verification of the adjointness axiom
$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
@@
To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$, we compute for arbitrary $n$ that
$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\l\left(Z_{+}\wedge\Delta[n]_{+}\right),Y\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(Z\times\Delta[n]\right)_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(Z\times\Delta[n],\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right).$
The second bijection uses Proposition 5.4, the third one uses Lemma 5.7, and the fourth one uses Lemma 2.8.
To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$, we compute for arbitrary $n$ that
$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\left(X\wedge\Delta[n]_{+}\right),Y\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{1}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
$=\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right).$
@@
The second bijection comes from Proposition 5.4, and the third bijection is a consequence of Lemma 2.23(2). ∎
5.5. Saturated classes
To verify the homotopy lifting-extension axiom, we adapt the proof of the homotopy lifting-extension axiom for simplicial sets in [GJ99, Chapter I, Proposition 5.2]. The proof for simplicial sets uses saturated classes of monomorphisms of simplicial sets, which are defined as follows.
Definition 5.9 ([GZ67, Chapter IV, Definition 2.1]).
A class $S$ of monomorphisms of simplicial sets is saturated if the following conditions are met.
  1. (1)
    All isomorphisms are contained in $S$.
  2. (2)
    In a pushout square
[…]
    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
  3. (3)
    In a section-retraction commutative diagram
[…]
    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
  4. (4)
    Given a sequence of composable maps
    $A_{1}\xrightarrow{m_{1}}A_{2}\xrightarrow{m_{2}}\cdots,$
    if $m_{t}$ is contained in $S$ for every positive integer $t$, then so is $A_{1}\rightarrow\operatorname{colim}(A_{n})$.
  5. (5)
    Given an indexing set $L$ and $m_{l}\colon A_{l}\rightarrow B_{l}$ for $l\in L$, if $m_{l}$ is contained in $S$ for every $l\in L$, then
    $\coprod\limits_{l\in L}m_{l}\colon\coprod\limits_{l\in L}A_{l}\rightarrow\coprod\limits_{l\in L}B_{l}$
    is contained in $S$.
As the intersection of saturated classes forms another saturated class, the following definition is well-posed.
Definition 5.10.
The saturation of a class $P$ of monomorphisms of simplicial sets is
$\langle P\rangle=\bigcap\limits_{\begin{subarray}{c}S\supset P\textrm{ and}\\ S\textrm{ is saturated }\end{subarray}}S.$
Saturated classes provide us with descriptions of the class of acyclic cofibrations of simplicial sets.
Proposition 5.11 ([GJ99, Chapter I, Proposition 4.2]).
Consider the monomorphism classes of simplicial sets
  1. (1)
    $P_{1}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{and }0\leq k\leq n\right\}$; and
  2. (2)
    $P_{2}=\left\{\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\rightarrow\Delta[1]\times\Delta[n]\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
Then, $\langle P_{1}\rangle=\langle P_{2}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$.
Attaching a disjoint basepoint yields a description of the class of acyclic cofibrations of pointed simplicial sets.
Corollary 5.12.
Consider the monomorphism classes of pointed simplicial sets
  1. (1)
    $(P_{1})_{+}=\left\{(j_{n,k})_{+}\colon\Lambda[n,k]_{+}\rightarrow\Delta[n]_{+}\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}$; and
  2. (2)
    $(P_{2})_{+}=\left\{\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
Then, $\langle(P_{1})_{+}\rangle=\langle(P_{2})_{+}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
We use Corollary 5.12 to prove a lemma that underpins our verification of the homotopy lifting-extension axiom.
Lemma 5.13.
For two pointed simplicial maps $i\colon A\rightarrow B$ and $j\colon X\rightarrow Y$, if both of them are monomorphisms, then the induced map
$\left(A\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B\wedge X\right)\rightarrow B\wedge Y$
is a monomorphism, which is also a weak homotopy equivalence if $i$ or $j$ is also a weak homotopy equivalence.
Proof.
If both $i$ and $j$ are monomorphisms, then so is the induced map. To prove the second part of the claim, the key idea is to write $P$ for the class of monomorphisms $A^{\prime}\rightarrow B^{\prime}$ of pointed simplicial sets such that
$\left(A^{\prime}\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B^{\prime}\wedge X\right)\rightarrow B^{\prime}\wedge Y$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. We show that $P$ contains the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. As $P$ is a saturated class, thus equal to its own saturation, Corollary 5.12 truncates our task to showing that
$(P_{2})_{+}\subset P.$
Thus, for $e\in\{0,1\}$ and $n\geq 0$, we need to show that
$\left(\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\wedge Y\right)\cup\left(\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge X\right)\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge Y$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. We use Lemma 5.7 to write down the commutative diagram
[…]
where the upper pair of vertical isomorphisms is an instance of Lemma 5.7 and
$\left(\Delta[1]_{+}\wedge\left(\left(\partial\Delta[n]_{+}\wedge Y\right)\cup\left(\Delta[n]_{+}\wedge X\right)\right)\right)\cup\left(\{e\}_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)\right)\rightarrow\Delta[1]_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$ because its geometric realization is the inclusion of a strong deformation retract in pointed topological spaces. Overall, we conclude that $(P_{2})_{+}\subset P$. ∎
5.6. Verification of the homotopy lifting-extension axiom
We are now ready to verify that the structure defined in Section 5.2 also satisfies the homotopy lifting-extension axiom.
Proposition 5.14.
If $i\colon A\rightarrow B$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $p\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, then the induced map
$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right)$
is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak homotopy equivalence.
Proof.
To prove that the induced map is a Kan fibration, we solve a lifting problem of the form
${{\Lambda[n,k]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{j_{n,k}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
The adjunction $(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon V$ from Lemma 2.8 yields the equivalent lifting problem
[…]
We apply Corollary 5.6 to infer that this new lifting problem is, in turn, equivalent to the lifting problem
${{\left(\Lambda[n,k]_{+}\wedge B\right)\underset{\left(\Lambda[n,k]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
By Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ as $\left(j_{n,k}\right)_{+}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved, so the induced map is a Kan fibration.
If $i$ or $p$ is also a weak homotopy equivalence, then we prove that the induced map is also a weak homotopy equivalence. To that end, we solve a lifting problem
${{\partial\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{i_{n}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
Another consecutive application of the adjunctions from Lemma 2.8 and Corollary 5.6 translates our lifting problem to the equivalent lifting problem
${{\left(\partial\Delta[n]_{+}\wedge B\right)\underset{\left(\partial\Delta[n]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
By Lemma 5.13, the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $\left(i_{n}\right)_{+}$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. Depending on whether $i$ or $p$ is also a weak homotopy equivalence, we have two cases to scrutinize.
  1. (1)
    If $i$ is also a weak homotopy equivalence, then, by Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
  2. (2)
    If $p$ is also a weak homotopy equivalence, then $p$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. As the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
In either case, the induced map is both a Kan fibration and a weak homotopy equivalence, as required. ∎
By Propositions 5.8 and 5.14, our definitions in Section 5.2 endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
@@ 5.6. Verification of the homotopy lifting-extension axiom
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
@@
Proof.
With our definition of function complexes, ${\mathcal{SS}\mathrm{ets}}_{1}$ is enriched over simplicial sets as ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is enriched over simplicial sets. By Propositions 5.8 and 5.14, our definitions satisfy the simplicial model category axioms. ∎
@@ 6. Rational homotopy theory of $1$-connected spaces
From now on, let $M$ be a multiplicative subset of $\mathbb{Z}$. We gather the background and auxiliary results from the rational homotopy theory of $1$-connected topological spaces that we need to recover the entirety of Quillen’s family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$. In particular, we characterize $M$-local spaces and $M$-local homotopy equivalences in terms of function complexes, which prepares us for the left Bousfield localization in Section 7. Even though we do not expect these characterizations to be novel to the rational homotopy theory expert, we were unable to find them in the literature, so we provide a full proof.
6.1. Localization of $1$-connected spaces
We first review some concepts from [FHT01, Chapter 9].

### v2

@@ 5.2. The simplicial model structure
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X))$.
With our definition of function complexes, ${\mathcal{SS}\mathrm{ets}}_{1}$ is enriched in simplicial sets by Example 5.3. The homotopy lifting-extension axiom also holds by Example 5.3, so it remains to verify the adjointness axiom.
@@
$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y))\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$
@@
The natural isomorphism $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y))$ is a consequence of Example 5.3. To prove that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$, we compute that
$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}((X\wedge Z_{+})\wedge\Delta[n]_{+},Y)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y))$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{1}}(X\wedge\Delta[n]_{+},E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$
$=\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y))),$
@@
where the first bijection comes from Example 5.3 and the second bijection follows from Proposition 2.19(2). ∎
Overall, our modification of Example 5.3 endows the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial structure.
@@
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X))$.
@@
Lastly, we note that, in [12], Burke establishes an analogous simplicial structure for reduced simplicial sets, that is, simplicial sets with a unique 0-simplex; Burke’s line of argument is different from ours, though.
@@ 6. Rational homotopy theory of $1$-connected spaces
From now on, let $M$ be a multiplicative subset of $\mathbb{Z}$. We gather the background and auxiliary results from the rational homotopy theory of $1$-connected topological spaces that we need to recover the entirety of Quillen’s family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$. In particular, we describe the $M$-local spaces and $M$-local homotopy equivalences in terms of function complexes, which prepares us for the left Bousfield localization in Section 7.
6.1. Rationalization of $1$-connected spaces
We give a review from [39, Chapter 2] and [22, Chapter 9].